OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (mertsj):

You could solve the first equation for y by subtracting 4x from both sides. Then you could substitute that answer for y into the second equation.

OpenStudy (anonymous):

I would just substitute the values in for x and y that you already have, but, for substitution you also need to solve for one of the equations. So, for the second one: x=8-2y. Then substitute that into the first equation: 4(8-2y)+y=11. Then solve. 32-8y+y=11. 32-7y=11, then -7y=-21, so y=3. And then substitute that value in for the y. So, 4x+3=11. 4x=8. x=2. Hope this helps

OpenStudy (anonymous):

2-3 ?

OpenStudy (mertsj):

Yes. (2,3) is the correct answer. Good for you!!